OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (1,3) m=3

OpenStudy (anonymous):

to find b (y intercept): y=mx+b 3=3(1) +b this is from (y=3 from the set 1,3)=(m=3)(x=1 from the set 1,3) 3=3+b 3-3=b b=0 The equation for the line in the form y = mx + b is: y=3x+0 to check y=3x+b 3=3(1)+0 3=3 it is correct